OpenStudy (anonymous):

Reduce 7y + 2y ^2 - 7 by 3 - 4y.

OpenStudy (marihelenh):

Same thing, subtract like terms. 2y^2-0=? 7y-(-4y)=7y+4y=? -7-3=?

OpenStudy (anonymous):

so -10?